1.4 Given the constraints (10)
A+B + C <= 24, B +C >=8 and A >= 0, B >= 0, C>=  0.
Maximize 24-A-B - C
A: amount of time spent on school work
B: amount of time spent on fun
C: amount of time spent on pay work
1
Expert's answer
2019-04-24T14:10:21-0400

Maximize 24-A-B - C

given the constraints A+B + C "\leq" 24, B +C "\geq8" and "A\geq0, B\geq0, C\geq0"

Then


"8\leq A+B+C\leq24"

"0\leq 24-A-B-C\leq16"

"24-A-B-C" has maximum value of 16 when

"A=0, B+C=8, B\geq0, C\geq0"